Can some help solve this story problem, please and thank you!
Tropical Pools sells an aboveground model for p dollars each.
The monthly revenue for this model is given by the formula:
R (p) -0.08p^2 + 300p (the formula)
a) Factor out the price on the right-hand side of the formula.
b) Write a formula D (p) for the monthly demand.
c) Find D (3000).

Answer by stanbon(75887)

The monthly revenue for this model is given by the formula:
R(p) = 0.08p^2 + 300p (the formula)
a) Factor out the price on the right-hand side of the formula.
R(p) = p(0.08p + 300)
--------------------------------
b) Write a formula D(p) for the monthly demand.
D(p) = 0.08p + 300
--------------------------------
c) Find D (3000).
D(3000) = 0.08*3000+300 = 540
====================================
Cheers,
Stan H.